complete the proof
given: 4x+6/2=9
prove: x=2
statements reasons
a.4x+6/2=9 a._______
b.___(4x+6/2)=2(9) b.multiplication prop.
c.4x+6=18 c._______
d. 4x+6-6=18-6 d._______
e. 4x= ____ e. substitution prop.
f.4x/4=____ f. division prop.
g.______ g. substitution prop.

Answer by jerryguo41(162)

a. Given
b. 2
c. 8x+6=18 Distribution
d. 8x+6-6=18-6 Subtraction Property
e. 8x=12 Division Prop.
f. x=12/8 Simplification
g. x= 1.5 or 1 and 1/2